# FlowDubber: Movie Dubbing with LLM-based Semantic-aware Learning and Flow Matching based Voice Enhancing

### v1

@@ Abstract.
Movie Dubbing aims to convert scripts into speeches that align with the given movie clip in both temporal and emotional aspects while preserving the vocal timbre of a given brief reference audio. Existing methods focus primarily on reducing the word error rate while ignoring the importance of lip-sync and acoustic quality. To address these issues, we propose a large language model (LLM) based flow matching architecture for dubbing, named FlowDubber, which achieves high-quality audio-visual sync and pronunciation by incorporating a large speech language model and dual contrastive aligning while achieving better acoustic quality via the proposed voice-enhanced flow matching than previous works. First, we introduce Qwen2.5 as the backbone of LLM to learn the in-context sequence from movie scripts and reference audio. Then, the proposed semantic-aware learning focuses on capturing LLM semantic knowledge at the phoneme level. Next, dual contrastive aligning (DCA) boosts mutual alignment with lip movement, reducing ambiguities where similar phonemes might be confused. Finally, the proposed Flow-based Voice Enhancing (FVE) improves acoustic quality in two aspects, which introduces an LLM-based acoustics flow matching guidance to strengthen clarity and uses affine style prior to enhance identity when recovering noise into mel-spectrograms via gradient vector field prediction. Extensive experiments demonstrate that our method outperforms several state-of-the-art methods on two primary benchmarks. The demos are available at https://galaxycong.github.io/LLM-Flow-Dubber/.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: arXiv Preprint; April, 2025; arXiv.org
@@ 1. Introduction
Movie Dubbing, also known as Visual Voice Cloning (V2C) (Chen et al., 2022), aims to generate a vivid speech from scripts using a specified timbre conditioned by a single short reference audio while ensuring strict audio-visual synchronization with lip movement from silent video, as shown in Figure 1. It attracts great attention in the multimedia community and promises significant potential in real-world applications such as film post-production and personal speech AIGC.
Figure 1. (a) Illustration of V2C task. (b) Illustration of the proposed FlowDubber. It brings a new level of high-quality lip-sync and pronunciation by incorporating a large speech language model and dual contrastive aligning while achieving better acoustic quality via voice-enhanced flow matching.
Previous dubbing works (Zhang et al., 2024, 2025a; Cong et al., 2024b, 2023; Chen et al., 2022) achieve significant progress in improving pronunciation and are dedicated to reducing the word error rate (WER) of generated speech. They can be mainly divided into two groups. Since the dubbing resources are limited in scale (copyright issues) and are always accompanied by background sounds or environmental noise, one class of methods (Zhang et al., 2024, 2025a) focuses primarily on leveraging external knowledge to improve pronunciation clarity by pre-training on clear large-scale text-to-speech corpus (Zen et al., 2019). For example, Speaker2Dubber (Zhang et al., 2024) proposes a two-stage dubbing architecture, which allows the model to first learn pronunciation via multi-task speaker pre-training on Libri-TTS 100 dataset and then optimize duration in stage two. Then, by pre-training on larger TTS corpus Libri-TTS 460 dataset, ProDubber (Zhang et al., 2025a) proposes another novel two-stage dubbing method based on Style-TTS2 model (Li et al., 2023), including prosody-enhanced pre-training and acoustic-disentangled prosody adapting. However, these pre-training methods rely too much on the TTS architecture (Ren et al., 2021; Li et al., 2023) and mainly adopt a Duration Predictor (DP) (Cong et al., 2024b) to produce rough duration without considering intrinsic relevance with lip motion, resulting in poor audio-visual sync.
@@
Except for the alignment issues mentioned above, the existing dubbing methods suffer from acoustic quality degradation, even in the advanced two-stage dubbing pre-training methods. For example, Speaker2Dubber (Zhang et al., 2024) freezes the text encoder in the second stage, which helps to maintain pronunciation. However, its use of a traditional FastSpeech2-based (Ren et al., 2021) transformer fails to handle the complex and diverse spectrum changes, leading to subpar acoustic quality. In addition, the authoritative acoustic quality measurement predictor UTMOS (Saeki et al., 2022) reveals that the acoustic quality of current dubbing methods still requires improvement.
Recent advances in speech tokenization (van den Oord et al., 2017; Du et al., 2024b; Mentzer et al., 2024; Ji et al., 2024a) have revolutionized TTS synthesis by bridging the fundamental gap between continuous speech signals and token-based large language models (LLM). Due to LLM demonstrating excellent capability in sequential modeling and contextual understanding, these LLMs-based speech synthesis models achieve human-level expressive and naturalness (Du et al., 2024b; Ye et al., 2025; Wang et al., 2025; Guo et al., 2024b). However, they are struggling to deal with dubbing task. Although some speed-controllable LLM speech models have been proposed, they still lack visual understanding capabilities, and the synthesized speech struggles to align with the lip motion changing in video. Besides, some speech models focus too much on the naturalness of speech, resulting in poor cloning ability, which makes it challenging to maintain speaker similarity in speech synthesis.
To address these issues, we propose an LLM-based flow matching architecture for dubbing, named FlowDubber, which guarantees high-quality audio-visual sync and pronunciation by incorporating a large speech language model and dual contrastive aligning while achieving state-of-the-art acoustic quality via the proposed voice-enhanced flow matching (as shown in Figure 1 (b)). Specifically, we first design an LLM-based Semantic-aware Learning (LLM-SL), which consists of pre-trained textual LLM Qwen2.5-0.5B to model the in-context sequence from movie scripts and reference audio and semantic-aware phoneme learning focuses on capturing the relevance between phoneme pronunciation unit and LLM semantic knowledge. Then, the proposed dual contrastive aligning (DCA) ensures mutual alignment between lip movement and phoneme sequence, reducing ambiguities where similar phonemes might be confused. Finally, we propose a novel Flow-based Voice Enhancing (FVE) module, which improves the acoustic quality from two sub-components: LLM-based acoustics flow matching guidance and style flow matching prediction. The LLM-based acoustics flow matching guidance focuses on improving the clarity of acoustics during recovering noise to mel-spectrograms by gradient vector field prediction conditioned on LLM.
@@
  • Extensive experimental results demonstrate the proposed FlowDubber  performs favorably against state-of-the-art models on two dubbing benchmark datasets.
@@ 2.1. Visual Voice Cloning
The V2C task (Chen et al., 2022) requires generating high-quality dubbing speech how a text might be said, but in step with the lip movements portrayed by video character, and in vocal style exemplified by reference audio (Kim et al., 2025; Sung-Bin et al., 2025; Choi et al., 2025, 2024; Ye and Shan, 2025; Zhang et al., 2025b; Zheng et al., 2025; Li et al., 2025). Some works focus primarily on improving the pronunciation clarity (Cong et al., 2023, 2024b; Zhao et al., 2024). For example, SOTA dubbing method ProDubber (Zhang et al., 2025a) and Speaker2Dubber (Zhang et al., 2024) propose a pre-training framework to learn clear pronunciation representation from a large-scale text-to-speech corpus (Zen et al., 2019). However, they over-rely on the TTS architecture and use an inaccurate duration predictor (Zhang et al., 2025a) to estimate the lip speaking time, without considering the intrinsic connection between visual movement and speech content. Besides, StyleDubber (Cong et al., 2024b) attempts to use time stretching to balance articulation and lip-sync. Although the overall length can remain consistent, this does not fundamentally achieve the audio-visual alignment in fine-grained lip-sync. In this work, we propose FlowDubber, a novel dubbing architecture that combines LLM-based semantic-aware learning with dual contrastive alignment to achieve high-quality lip synchronization, and the proposed flow-matching enhancing mechanism delivers better acoustic quality than existing dubbing methods.
Figure 2. Overall framework of FlowDubber. It consists of LLM-based Semantic-aware Learning (LLM-SL), lip-phoneme Dual Contrastive Aligning (DCA), and Flow-based Voice Enhancing (FVE). Specifically, the LLM-SL includes Qwen2.5-0.5B speech language model and semantic-aware phoneme learning to keep pronunciation while aligning with DCA. The FVE consists of style flow matching prediction and LLM-based acoustics flow matching guidance to improve the acoustics quality.
@@ 2.2. Large Language Model and Speech Codec
The remarkable success of Large Language Models (LLMs) (Brown et al., 2020; Guo et al., 2025; Yang et al., 2024b) and the autoregressive (AR) model brings significant advancements in the field of speech synthesis. VALL-E (Wang et al., 2023a) first converts speech into neural codec tokens and treats the speech synthesis as a next-token prediction task. Subsequently, extensive research focuses on speech codecs and LLM-based speech generators to improve the synthesis performance. For example, DAC (Kumar et al., 2023) adopts the residual vector quantization and the multi-scale STFT discriminators to obtain higher-quality discrete speech tokens. Wavtokenizer (Ji et al., 2024b) and X-codec (Ye et al., 2024) further improved the efficiency of codec and addressed the semantic shortcomings of previous codes. Besides, LLM-based speech synthesis systems combine the AR model with other components (Anastassiou et al., 2024; Chen et al., 2024a) or rely on continuous acoustic features (Meng et al., 2024; Zhu et al., 2024) to achieve better performance. Recently, Llasa (Ye et al., 2025) investigated the effects of training-time inference-time scaling in LLM-based speech synthesis. However, they still lack visual understanding capability, and the generated speech struggles to align with the lip movement. In this paper, we propose powerful dubbing model that can achieve the best lip-sync and inherit the acoustic knowledge of LLM via flow-matching learning and phoneme learning, achieving advanced results against all dubbing methods.
@@ 2.3. Speech Synthesis and Flow Matching
Flow Matching (Lipman et al., 2022) is a simulation-free approach to training continuous normalizing flow (Chen et al., 2018) models, capable of modeling arbitrary probability paths and capturing the trajectories represented by diffusion processes (Song et al., 2021). Due to the high quality and faster speed, flow matching has attracted significant attention in speech generation (Guo et al., 2024a; Kim et al., 2023; Le et al., 2023). For example, Matcha-TTS (Mehta et al., 2024) adopts the optimal transport conditional flow matching in single speaker TTS synthesis and Stable-VC (Yao et al., 2024) adopts it in voice conversion field to improve fidelity. F5-TTS (Chen et al., 2024b) is another powerful TTS model to reconstruct high-quality mel-spectrograms by flow matching. Then, CosyVoice 2.0 (Du et al., 2024a, b) has further proven its superior performance by combining flow matching with LLM. However, these methods are not suited to V2C dubbing task due to they inability to perceive proper pause in step with lip motion. Recently, EmoDub (Cong et al., 2024a) introduces classifier guidance in flow matching to control emotions via input labels and intensity. In contrast, after integrating semantic-aware phoneme learning and lip-motion aligning, we focus on refining the flow-matching generation process to ensure clarity by introducing semantic knowledge from LLM via classifier-free guidance.
@@ 3.1. Overview
where the $V_{s}$ represents the given a silent video clip, $W_{r}$ is a reference waveform used for voice cloning, and $T_{c}$ is current piece of text to convey speech content. The goal of FlowDubber  is to generate a piece of high-quality speech $\hat{Y}$ that guarantees precise lip-sync with silent video, high speaker similarity, and clear pronunciation. The main architecture of the proposed model is shown in Figure 2. Specifically, we introduce pre-trained textual LLM Qwen2.5-0.5B as the backbone of the speech language model to model the in-context sequence from movie scripts and reference audio by discretizing them. Then, the semantic knowledge of speech tokens is adapted to the phoneme level by semantic-aware phoneme learning. Next, the proposed Dual Contrastive Aligning (DCA) ensures the cross model alignment between lip-motion and phoneme level information from LLM. Finally, Flow-based Voice Enhancement (FVE) enhances the fused information from two aspects: Style Flow Matching Prediction aims to keep the speaker similarity and LLM-based Acoustics Flow Matching Guidance focuses on improving the acoustics clarity and suppressing noise. We detail each module below.
@@ 3.2. LLM-based Semantic-aware Learning
Different from the previous dubbing works (Cong et al., 2023; Zhao et al., 2024), we introduce LLM-based semantic-aware learning to capture the phoneme level pronunciation via the powerful in-context learning capabilities of LLM (Qwen2.5-0.5B) between text token in movie script and semantic and identity token in reference audio.
Speech Tokenization. This module aims to transform the speech signal of reference audio $R_{a}$ into a sequence of semantic tokens ${h}_{q}$. It first utilizes a pre-trained self-supervised learning (SSL) model, wav2vec 2.0 (Baevski et al., 2020), to translate speech signals into a semantic embedding sequence. Then, the semantic encoder $S_{encoder}(\cdot)$, constructed with 12 ConvNeXt (Liu et al., 2022) blocks and 2 downsampling blocks, is employed to process and down-sample the sequence further into an encoding sequence $h$:
@@
where the output ${H}_{q}$ represents semantic tokens from $h$ by Vector Quantization $\text{VQ}(\cdot)$ layers. Specifically, the $\text{VQ}(\cdot)$ adopts factorized codes manner and has an 8192 codebook size and 8 codebook dimensions. Different from CosyVocie 2.0 (Du et al., 2024b), we also extract global tokens ${G}_{q}$ from reference audio’s mel-spectrogram by Finite Scalar Quantization (FSQ) (Mentzer et al., 2024) to keep the speaker characteristics (Wang et al., 2025).
Speech Language Model. Inspired by LLM successes, we employ the pre-trained textual LLM Qwen2.5-0.5B (Yang et al., 2024b) as the backbone of the speech language model. Specifically, we formulate GPT (Radford et al., 2019) architecture as the next-token prediction paradigm, which adopts a decoder-only autoregressive transformer architecture:
@@
Phoneme Level Semantic-aware Module. Compared with zero-shot TTS, movie dubbing must be strictly matched with lip movements from silent video to achieve audio-visual synchronization. The proposed phoneme-level semantic-aware module aims to capture the semantic knowledge from the speech language model at the phoneme level, which helps preserve pronunciation and enables fine-grained alignment between phoneme unit and lip motion sequence. Specifically, the phoneme-level semantic-aware module consists of cross-modal transformers $\hat{Z}^{[i]}_{S\rightarrow P}$ to calculate the relevance between textual phoneme embedding and LLM speech knowledge, which can be formulate as:
(4) $\begin{split}\hat{Z}^{[i]}_{S\rightarrow P}=&\;\text{LLM}^{[i],\text{mul}}_{S\rightarrow P}(\text{LN}(Z^{[i-1]}_{S\rightarrow P}),\text{LN}(Z^{[0]}_{S}))+\text{LN}(Z^{[i-1]}_{S\rightarrow P}),\\ Z^{[i]}_{S\rightarrow P}=&\;f_{\theta^{[i]}_{S\rightarrow P}}(\text{LN}(\hat{Z}^{[i]}_{S\rightarrow P})+\text{LN}(\hat{Z}^{[i]}_{S\rightarrow P}),\end{split}$
@@
[…]

### v2

@@ Abstract.
Movie Dubbing aims to convert scripts into speeches that align with the given movie clip in both temporal and emotional aspects while preserving the vocal timbre of a given brief reference audio. Existing methods focus primarily on reducing the word error rate while ignoring the importance of lip-sync and acoustic quality. To address these issues, we propose a novel dubbing architecture based on Large Language Model (LLM) and Conditional Flow Matching (CFM), named FlowDubber, which achieves high-quality audio-visual sync and pronunciation by incorporating a large speech language model with dual contrastive alignment while improving acoustic quality via Flow-based Voice Enhancing (FVE). First, we introduce Qwen2.5 as the backbone of large speech language model to learn the in-context sequence from movie scripts and reference audio. Second, the proposed semantic-aware learning focuses on capturing LLM semantic knowledge at the phoneme level, which facilitates mutual alignment with lip movement from silent video via Dual Contrastive Alignment (DCA). Third, the FVE introduces an LLM-based acoustics flow matching guidance to strengthen clarity by decoupling Classifier-Free Guidance (CFG) enhancement. Extensive experiments demonstrate that our method outperforms several state-of-the-art methods on two primary benchmarks. The demos are available at https://galaxycong.github.io/LLM-Flow-Dubber/.
@@
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754734isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Phonology / morphologyccs: Computing methodologies Computer vision
@@ 1. Introduction
Movie Dubbing, also known as Visual Voice Cloning (V2C) (Chen et al., 2022), aims to generate a vivid speech from scripts using a specified timbre conditioned by a single short reference audio while ensuring strict audio-visual synchronization with lip movement from silent video, as shown in Figure 1(a). It attracts great attention in the multimedia community and promises significant potential in real-world applications such as film post-production and personal speech AIGC.
Figure 1. (a) V2C task. (b) Unlike dubbing with duration predictor from TTS, FlowDubber incorporates a large language model (LLM) and voice-enhanced flow matching to generate high-quality dubbing while ensuring lip-sync.
Previous dubbing works (Zhang et al., 2024b, 2025a; Cong et al., 2024b, 2023; Chen et al., 2022) achieve significant progress in improving pronunciation and are dedicated to reducing the word error rate (WER) of generated speech. They can be mainly divided into two groups. Since the dubbing resources are limited in scale (copyright issues) and are always accompanied by background sounds or environmental noise, one class of methods (Zhang et al., 2024b, 2025a) focuses primarily on leveraging external knowledge to improve pronunciation clarity by pre-training on clear, large-scale text-to-speech corpus (Zen et al., 2019). For example, Speaker2Dubber (Zhang et al., 2024b) proposes a two-stage dubbing architecture, which allows the model to first learn pronunciation via multi-task speaker pre-training on Libri-TTS 100 dataset and then optimize duration in stage two. Then, by pre-training on larger TTS corpus Libri-TTS 460 dataset, ProDubber (Zhang et al., 2025a) proposes another novel two-stage dubbing method based on the Style-TTS2 model (Li et al., 2023), including prosody-enhanced pre-training and acoustic-disentangled prosody adapting. However, these pre-training methods rely too much on the TTS architecture (Ren et al., 2021; Li et al., 2023) and mainly adopt a Duration Predictor (DP) (Cong et al., 2024b) to produce rough duration without considering intrinsic relevance with lip motion, resulting in poor audio-visual sync.
@@
Except for the alignment issues mentioned above, the existing dubbing methods suffer from acoustic quality degradation, even in the advanced two-stage dubbing pre-training methods. For example, Speaker2Dubber (Zhang et al., 2024b) freezes the text encoder in the second stage, which helps to maintain pronunciation. However, its use of a traditional FastSpeech2-based (Ren et al., 2021) transformer fails to handle the complex and diverse spectrum changes, leading to subpar acoustic quality. In addition, the acoustic quality measurement predictor UTMOS (Saeki et al., 2022) reveals that the acoustic quality of current dubbing methods still requires improvement.
Recent advances in speech tokenization (van den Oord et al., 2017; Du et al., 2024b; Mentzer et al., 2024; Ji et al., 2024a) have revolutionized TTS synthesis by bridging the fundamental gap between continuous speech signals and token-based large language models (LLM). Due to LLM demonstrating excellent capability in sequential modeling and contextual understanding, these LLM-based speech synthesis models achieve human-level expressive and naturalness (Du et al., 2024b; Ye et al., 2025b; Wang et al., 2025; Guo et al., 2024b). However, they are struggling to deal with the dubbing task. Although some speed-controllable LLM speech models have been proposed, they still lack visual understanding capabilities, and the synthesized speech struggles to align with the lip motion changing in video.
To address these issues, we propose an LLM-based flow matching architecture for dubbing, named FlowDubber (as shown in Figure 1 (b)), which incorporates a large speech language model and dual contrastive alignment to ensure audio-visual sync and pronunciation, while achieving better acoustic quality via voice-enhanced flow matching than the state-of-the-art method. Specifically, we first introduce an LLM-based Semantic-aware Learning (LLM-SL), which leverages pre-trained LLM Qwen2.5-0.5B to model the in-context sequence from movie scripts (text) and reference audio (reference token including ref. semantic and ref. global token). Then, the proposed semantic-aware phoneme learning captures the connection between phoneme-level pronunciations and LLM-derived semantics, making them well-suited for integration into the Dual Contrastive Aligning (DCA) module. Next, the DCA is designed to perform mutual alignment between lip movement and phoneme sequence to ensure lip-sync. Finally, we propose a novel Flow-based Voice Enhancing (FVE) module, which improves the acoustic quality from two sub-components: LLM-based acoustics flow matching guidance and style flow matching prediction. The key part is LLM-based acoustics flow matching guidance, which focuses on improving clarity during recovering noise to mel-spectrograms by decoupling Classifier-Free Guidance (CFG) enhancement.
@@
  • Extensive experimental results demonstrate the proposed FlowDubber  performs favorably against state-of-the-art models on two popular dubbing benchmark datasets.
@@ 2.1. Visual Voice Cloning
With the rapid development of deep learning (Tu et al., 2024; Liu et al., 2023; Zhang et al., 2024c; Ye et al., 2022; Zhang et al., 2024a; Li et al., 2025a), V2C (Chen et al., 2022) has attracted great interest in the multimedia community (Choi et al., 2025, 2024; Tang et al., 2024, 2025; Zhang et al., 2024d; Cui et al., 2024; Ye et al., 2023, 2025a). It requires determining how a text should be spoken, in sync with the lip movements in silent video and in the vocal style of reference audio (Kim et al., 2025; Sung-Bin et al., 2025; Lee et al., 2023a; Ye and Shan, 2025; Zhang et al., 2025b; Zheng et al., 2025; Li et al., 2025b). Some V2C works focus primarily on improving the pronunciation clarity (Cong et al., 2023, 2024b; Zhao et al., 2024). For example, SOTA dubbing method ProDubber (Zhang et al., 2025a) and Speak2Dub (Zhang et al., 2024b) propose a two-stage framework to learn clear pronunciation by pre-training from large-scale TTS corpus (Zen et al., 2019). However, they over-rely on the TTS architecture and use an inaccurate duration predictor (Zhang et al., 2025a) to estimate the lip speaking time, without considering the intrinsic audio-visual alignment. Besides, StyleDubber (Cong et al., 2024b) uses time stretching in the duration predictor. Although the overall length of the dubbing can be consistent, it does not fundamentally capture fine-grained lip-sync with the video. In this work, we propose FlowDubber, a novel dubbing architecture that combines LLM-based semantic-aware learning with dual contrastive alignment to achieve high-quality lip synchronization, and flow-matching enhancing mechanism is designed to achieve better acoustic quality than existing dubbing methods.
Figure 2. Overall framework of FlowDubber. It consists of LLM-based Semantic-aware Learning (LLM-SL), Dual Contrastive Aligning (DCA), and Flow-based Voice Enhancing (FVE). Specifically, the LLM-SL includes Qwen2.5-0.5B speech language model and semantic-aware phoneme learning to keep pronunciation while ensuring lip-sync by DCA. The FVE is equipped with LLM-based Acoustics Flow Matching Guidance and Style Flow Matching Prediction to improve clarity and similarity.
@@ 2.2. Large Language Model and Speech Codec
The remarkable success of Large Language Models (LLMs) (Brown et al., 2020; Guo et al., 2025; Yang et al., 2024) and the autoregressive (AR) model brings significant advancements in the field of speech synthesis. VALL-E (Wang et al., 2023a) first converts speech into neural codec tokens and treats the speech synthesis as a next-token prediction task. Subsequently, extensive research focuses on speech codecs and LLM-based speech generators to improve the synthesis performance. For example, DAC (Kumar et al., 2023) adopts the residual vector quantization and the multi-scale STFT discriminators to obtain higher-quality discrete speech tokens. Wavtokenizer (Ji et al., 2024b) and X-codec (Ye et al., 2024) further improved the efficiency of codec and addressed the semantic shortcomings of previous codes. Besides, LLM-based speech synthesis systems combine the AR model with other components (Anastassiou et al., 2024; Chen et al., 2024a) or rely on continuous acoustic features (Meng et al., 2024; Zhu et al., 2024) to achieve better performance. Recently, Llasa (Ye et al., 2025b) investigated the effects of training-time inference-time scaling in LLM-based speech synthesis. However, they still lack visual understanding capability, and the generated speech struggles to align with the lip movement. In this paper, we propose an effective dubbing model that can achieve high-quality audio-visual alignment and inherit the acoustic knowledge from LLM via Semantic-aware Phoneme Learning and LLM-based Acoustics Flow Matching Guidance.
@@ 2.3. Speech Synthesis and Flow Matching
Flow Matching (Lipman et al., 2022) is a simulation-free approach to training continuous normalizing flow (Chen et al., 2018) models, capable of modeling arbitrary probability paths and capturing the trajectories represented by diffusion processes (Song et al., 2021). Due to the high quality and faster speed, flow matching has attracted significant attention in speech generation (Guo et al., 2024a; Kim et al., 2023; Le et al., 2023; Zhang et al., 2025c). Matcha-TTS (Mehta et al., 2024) adopts the optimal transport conditional flow matching in single speaker TTS synthesis, and Stable-VC (Yao et al., 2024) adopts it in voice conversion field to improve fidelity. F5-TTS (Chen et al., 2024b) is another powerful TTS model to reconstruct high-quality mel-spectrograms by flow matching. Then, CosyVoice 2.0 (Du et al., 2024a, b) has further proven its superior performance by combining flow matching with LLM. However, these methods are not suited to the V2C dubbing task due to their inability to perceive proper pauses in step with lip motion. Recently, EmoDub (Cong et al., 2024a) introduces classifier guidance in flow matching to control emotions via input labels and intensity. In contrast, after integrating semantic-aware phoneme learning and lip-motion aligning, we focus on refining the flow-matching generation process to ensure clarity by introducing semantic knowledge from LLM via classifier-free guidance.
@@ 3.1. Overview
where the $V_{s}$ represents the given silent video clip, $W_{r}$ is a reference waveform used for voice cloning, and $T_{c}$ is current piece of text to convey speech content. The goal of FlowDubber  is to generate a piece of high-quality speech $\hat{Y}$ that guarantees precise lip-sync with silent video, high speaker similarity, and clear pronunciation. The main architecture of the proposed model is shown in Figure 2. Specifically, we introduce pre-trained textual LLM Qwen2.5-0.5B as the backbone of the speech language model to learn the in-context sequence from movie scripts and reference audio by discretizing them. Then, the semantic knowledge of speech tokens is adapted to the phoneme level by semantic-aware phoneme learning. Next, the proposed Dual Contrastive Aligning (DCA) ensures the mutual alignment between lip-motion and phoneme-level information from LLM. Finally, Flow-based Voice Enhancing (FVE) aims to maintain the speaker’s similarity and improve the clarity by an LLM-based Acoustics Flow Matching Guidance. We detail each module below.
@@ 3.2. LLM-based Semantic-aware Learning
Different from the previous dubbing works (Cong et al., 2023; Zhao et al., 2024), we introduce LLM-based semantic-aware learning to capture the phoneme-level pronunciation via the powerful in-context learning capabilities of LLM (Qwen2.5-0.5B) between text tokens in movie script and semantic and global tokens in reference audio.
Speech Tokenization. This module aims to transform the speech signal of reference audio $R_{a}$ into a sequence of semantic tokens ${h}_{q}$, following Spark-TTS (Wang et al., 2025). It first utilizes a pre-trained self-supervised learning (SSL) model, wav2vec 2.0 (Baevski et al., 2020), to translate speech signals into a semantic embedding sequence. Then, the semantic encoder $S_{encoder}(\cdot)$, constructed with 12 ConvNeXt (Liu et al., 2022) blocks and 2 downsampling blocks, is employed to process and down-sample the sequence further into an encoding sequence $h$:
@@
where the output ${H}_{q}$ represents semantic tokens. $\text{VQ}(\cdot)$ adopts a factorized code structure with a codebook size of 8192 and 8 codebook dimensions. ${G}_{q}$ denotes the global tokens by Finite Scalar Quantization (FSQ), following Spark-TTS.  (Wang et al., 2025).
Speech Language Model. Inspired by LLM successes, we employ the pre-trained Qwen2.5-0.5B (Wang et al., 2025) as the backbone of the speech language model. Specifically, we formulate the GPT (Radford et al., 2019) architecture as the next-token prediction paradigm, which adopts a decoder-only autoregressive transformer architecture:
@@
Phoneme Level Semantic-aware Module. Compared with zero-shot TTS, movie dubbing must be strictly matched with lip movements from silent video to achieve audio-visual synchronization. The proposed phoneme-level semantic-aware module aims to capture the semantic knowledge from the speech language model at the phoneme level, which helps preserve pronunciation and enables fine-grained alignment between phoneme unit and lip motion sequence. Specifically, the phoneme-level semantic-aware module consists of cross-modal transformers $\hat{Z}^{[i]}_{S\rightarrow P}$ to calculate the relevance between textual phoneme embedding and LLM speech knowledge, which can be formulated as:
[…]
@@
where $\text{G2P}(\cdot)$ denotes the grapheme-to-phoneme to convert raw text $T_{c}$ to a phoneme sequence, then the phoneme encoder $E_{pho}(\cdot)$ is used to obtain textual phoneme embedding. The $S_{llm}$ indicates the mapping speech feature from LLM token sequence $o_{1:N_{o}}$ by codec decoder (Wang et al., 2025). In this case, the $S_{llm}$ is used as key and value, and the textual phoneme embedding is used as query. Finally, we denote the last layer output of cross modal transformer as ${LLM}_{p}\in\mathbb{R}^{{l_{p}}\times{d_{m}}}$, which represents the phoneme-level semantic feature from LLM. The $l_{p}$ denotes the length of phoneme sequences and $d_{m}$ is the embedding size.